# TWISTED CONJUGACY IN DIHEDRAL ARTIN GROUPS I: TORUS KNOT GROUPS

### v3

@@ INTRODUCTION
Definition 4.1.
Let $G=\langle X\rangle$. For fixed $k\in\mathbb{N}$, the $k$-simultaneous conjugacy problem takes as input two $k$-tuples $(y_{1},\dots,y_{k}),(z_{1},\dots,z_{k})$, where each $y_{i},z_{i}\in X^{\ast}$, and decides whether there exists an element $g\in G$ such that $g^{-1}y_{i}g=_{G}z_{i}$ for all $i=1,\dots,k$. We say this is an effective solution if such an algorithm also produces a conjugator $g\in G$.
This decision problem is also known as conjugacy for finite lists, and some results on this decision problem can be found in [15, 16]. Note there exists finitely presented groups with solvable conjugacy problem and unsolvable SCP [4]. The following result will be useful later.
Theorem 4.2.
@@
Let $y_{i}=(y_{i_{1}},y_{i_{2}})$, $z_{i}=(z_{i_{1}},z_{i_{2}})$ for some $1\leq i\leq k$. By [13, Proposition 3.1], checking if $y_{i}$ and $z_{i}$ are conjugate is equivalent to checking that $y_{i_{2}}=z_{i_{2}}$, and that $y_{i_{1}}$ is a cyclic conjugate of $z_{i_{1}}$. In particular, the SCP for $G(m)$ is equivalent to solving the SCP for $\mathbb{Z}_{2}\ast\mathbb{Z}_{m}$ when $m$ is odd, or $\mathbb{Z}_{m}\ast\mathbb{Z}$ when $m$ is even. In both cases, these are hyperbolic groups, which have solvable SCP by [4, Theorem A]. Moreover, this solution is effective. ∎
@@
Proposition 4.3.
Let $G(m)=\langle X\rangle$ be an odd dihedral Artin group. Let $\varphi=\psi\circ\phi\in\mathrm{Aut}(G(m))$, where $\psi\in\mathrm{Inn}(G(m))$, $\phi\in\mathrm{Out}(G(m))$ such that for all $w\in G(m)$, $\psi(w)=g^{-1}wg$ for some $g\in G(m)$. Then there exists an algorithm which takes as input $\varphi\in\mathrm{Aut}(G(m))$, and determines the element $g\in G(m)$ defined by $\psi\in\mathrm{Inn}(G(m))$.
Moreover, the algorithm runs in time $\mathcal{O}(2\mu)$, where $\mu=4\cdot\mathrm{max}\{1,l(\varphi(x)),l(\varphi(y))\}$. In particular, the algorithm runs in linear time based on the generators $x,y\in X$ and the automorphism $\varphi\in\mathrm{Aut}(G(m))$ given.
Proof.
By 2.8, we have two cases to consider. To check if $\phi\in\mathrm{Out}(G(m))$ is trivial, we need to check if there exists $g\in G(m)$ such that $g^{-1}xg=_{G(m)}\varphi(x)$ and $g^{-1}yg=_{G(m)}\varphi(y)$. This is an example of the SCP in $G(m)$, for which we can determine the element $g\in G(m)$ by 4.2. Otherwise, suppose $\phi\in\mathrm{Out}(G(m))$ is of the form in Eq. 3. In this case, we need to check if there exists $g\in G(m)$ such that $g^{-1}\phi(x)g=g^{-1}x^{-1}g=_{G(m)}\varphi(x)$, and $g^{-1}\phi(y)g=g^{-1}y^{-1}g=_{G(m)}\varphi(y)$. Again the element $g\in G(m)$ can be determined using the SCP.
For complexity, the inputs of the SCP algorithm are $(x^{\varepsilon_{x}},y^{\varepsilon_{y}}),(\varphi(x),\varphi(y))$, where $\varepsilon_{x},\varepsilon_{y}=\pm 1$. By [5, Theorem 1], the algorithm has complexity as stated, noting that we have to account for all 4 cases for the different values of $\varepsilon_{x}$ and $\varepsilon_{y}$. ∎
@@
Proof.
Let $u,v\in X^{\ast}$, and let $\varphi=\psi\circ\phi\in\mathrm{Aut}(G(m))$ as in 4.3 be our inputs. We want to decide whether there exists $w\in X^{\ast}$ such that $v=_{G(m)}\varphi(w)^{-1}uw$. We will use 2.3 to reduce the problem to twisted conjugacy with respect to outer automorphisms. Here $\psi\in\mathrm{Inn}(G(m))$, i.e. $\psi(w)=_{G(m)}g^{-1}wg$ for some $g\in G(m)$, which can be determined in linear time by 4.3. We need to check two cases:
  1. 1.
    $u\sim_{\psi}v$: Here $v=_{G(m)}\psi(w)^{-1}uw=g^{-1}w^{-1}guw$. Rearranging gives $gv=_{G(m)}w^{-1}(gu)w$, and so this case reduces to solving the conjugacy problem with respect to $(gu,gv)$. Since $g$ is known, this is solvable in linear time by [13, Proposition 3.1].
  2. 2.
    $u\sim_{\psi\circ\phi}v$, where $\phi\in\mathrm{Out}(G(m))$ is of the form in Eq. 3: Here $v=_{G}\psi(\phi(w))^{-1}uw=g^{-1}\phi(w)^{-1}guw$. Rearranging gives $gv=_{G}\phi(w)^{-1}(gu)w$, and so this case reduces to solving the twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$ with respect to $\phi\in\mathrm{Out}(G(m))$. This is solvable in linear time by 3.21.
CONJUGACY PROBLEM IN EXTENSIONS OF $G(M)$
@@
Definition 5.2.
Let $A\leq\mathrm{Aut}(G)$ for a group $G=\langle X\rangle$. The orbit decidability problem for $A$, denoted OD(A), takes as input two words $u,v\in X^{\ast}$, and decides whether there exists $\phi\in A$ such that $v\sim\phi(u)$.
When $m$ is odd, we recall that $|\mathrm{Out}(G(m))|=2$ by 2.8, so we can apply a similar technique as [12, Theorem 5.1] to prove Section 1.
Proof of Section 1.
Let $\varphi_{1},\dots\varphi_{s}\in\mathrm{Aut}(G(m))$ be given, and consider $A=\langle\varphi_{1},\dots\varphi_{s}\rangle\leq\mathrm{Aut}(G(m))$. For each $i=1,\dots,s$, compute $w_{i}\in G(m)$ such that $\varphi_{i}=\gamma_{w_{i}}\phi_{i}$, where $\gamma_{w_{i}}\in\mathrm{Inn}(G(m))$ and $\phi_{i}\in\mathrm{Out}(G(m))$ (see 4.3). Given two words $u,v\in X^{\ast}$, we want to decide whether $v\sim\varphi(u)$ for some $\varphi\in A$.
If $\phi_{i}$ is trivial, for every $i$, then $A\leq\mathrm{Inn}(G(m))$, and so the set $\{\varphi(u)\mid\varphi\in A\}$ is a collection of conjugates of $u$. Here our problem reduces to deciding if $v$ is conjugate to $u$, which is decidable by [13, Proposition 3.1]. Otherwise, the set $\{\varphi(u)\mid\varphi\in A\}$ is a collection of conjugates of $u$ and $\phi(u)$, where $\phi\in\mathrm{Out}(G(m))$ as in Eq. 3. If $m$ is odd, we need to decide whether $v$ is conjugate to either $u$ or $\phi(u)$, which is decidable by two applications of [13, Proposition 3.1]. ∎
Our final result is then immediate from Section 1, Section 1 and 5.1.
@@ Acknowledgements
The author would like to thank Laura Ciobanu and Fujii Michihiko for helpful discussions. They would also like to thank Andrew Duncan and Alessandro Sisto for their comments.
@@ Appendix A Geodesic normal forms
We note that for $\epsilon=\pm 1$, then any pairs $\left(y^{\epsilon(k+1)},y^{-\epsilon(k+1)}\right)$ can be rewritten as $\left(y^{-\epsilon k},y^{\epsilon k}\right)$, using the rewrite rule
@@
This covers all cases, and we obtain a geodesic representative in linear time. ∎
@@
Table 2: Geodesic normal forms

### v5

@@ 4. Algorithm for the twisted conjugacy problem
{defi}
Let $G=\langle X\rangle$. For fixed $k\in\mathbb{N}$, the $k$-simultaneous conjugacy problem takes as input two $k$-tuples $(y_{1},\dots,y_{k}),(z_{1},\dots,z_{k})$, where each $y_{i},z_{i}\in X^{\ast}$, and decides whether there exists an element $g\in G$ such that $g^{-1}y_{i}g=_{G}z_{i}$ for all $i=1,\dots,k$. We say this is an effective solution if such an algorithm also produces a conjugator $g\in G$. This decision problem is also known as conjugacy for finite lists, and some results on this decision problem can be found in [15, 16]. Note there exists finitely presented groups with solvable conjugacy problem and unsolvable SCP [4]. The following result will be useful later.
Theorem 21.
@@
Let $y_{i}=(y_{i,1},y_{i,2})$, $z_{i}=(z_{i,1},z_{i,2})$ for some $1\leq i\leq k$. By [13, Proposition 3.1], checking if $y_{i}$ and $z_{i}$ are conjugate is equivalent to checking that $y_{i,2}=z_{i,2}$, and that $y_{i,1}$ is a cyclic conjugate of $z_{i,1}$. In particular, the SCP for $G(m)$ is equivalent to solving the SCP for $\mathbb{Z}_{2}\ast\mathbb{Z}_{m}$ when $m$ is odd, or $\mathbb{Z}_{m}\ast\mathbb{Z}$ when $m$ is even. In both cases, these are hyperbolic groups, which have solvable SCP by [4, Theorem A]. Moreover, this solution is effective.
@@
Proposition 22.
Let $G(m)=\langle X\rangle$ be an odd dihedral Artin group. Let $\psi=\iota_{g}\phi\in\mathrm{Aut}(G(m))$. Then there exists an algorithm which takes as input $\psi\in\mathrm{Aut}(G(m))$, and determines the element $g\in G(m)$ defined by $\iota_{g}\in\mathrm{Inn}(G(m))$.
Moreover, the algorithm runs in time $\mathcal{O}(2\mu)$, where $\mu=4\cdot\mathrm{max}\{1,l(\psi(x)),l(\psi(y))\}$. In particular, the algorithm runs in linear time based on the generators $x,y\in X$ and the automorphism $\psi\in\mathrm{Aut}(G(m))$ given.
Proof 4.2.
By Theorem 4, we have two cases to consider. To check if $\phi\in[\psi]$ is trivial, we need to check if there exists $g\in G(m)$ such that $g^{-1}xg=_{G(m)}\psi(x)$ and $g^{-1}yg=_{G(m)}\psi(y)$. This is an example of the SCP in $G(m)$, for which we can determine the element $g\in G(m)$ by Theorem 21. Otherwise, suppose $\phi\in\mathrm{Aut}(G(m))$ is of the form in Eq. 3. In this case, we need to check if there exists $g\in G(m)$ such that $g^{-1}\phi(x)g=g^{-1}x^{-1}g=_{G(m)}\psi(x)$, and $g^{-1}\phi(y)g=g^{-1}y^{-1}g=_{G(m)}\psi(y)$. Again the element $g\in G(m)$ can be determined using the SCP.
For complexity, the inputs of the SCP algorithm are $(x^{\varepsilon_{x}},y^{\varepsilon_{y}}),(\psi(x),\psi(y))$, where $\varepsilon_{x},\varepsilon_{y}\in\{\pm 1\}$. By [5, Theorem 1], the algorithm has complexity as stated, noting that we have to account for all 4 cases for the different values of $\varepsilon_{x}$ and $\varepsilon_{y}$.
@@
Proof 4.3.
Let $u,v\in X^{\ast}$, and let $\psi=\iota_{g}\phi\in\mathrm{Aut}(G(m))$ be our inputs. We want to decide whether there exists $w\in X^{\ast}$ such that $v=_{G(m)}\psi(w)^{-1}uw$. We will use Remark 1 to reduce the problem to twisted conjugacy with respect to outer automorphisms. Recall $\iota_{g}\in\mathrm{Inn}(G(m))$, i.e. $\iota_{g}(w)=_{G(m)}g^{-1}wg$ for some $g\in G(m)$, which can be determined in linear time by Proposition 22. We need to check two cases:
  1. (1)
    $u\sim_{\iota_{g}}v$: Here $v=_{G(m)}\iota_{g}(w)^{-1}uw=g^{-1}w^{-1}guw$. Rearranging gives $gv=_{G(m)}w^{-1}(gu)w$, and so this case reduces to solving the conjugacy problem with respect to $(gu,gv)$. Since $g$ is known, this is solvable in linear time by [13, Proposition 3.1].
  2. (2)
    $u\sim_{\iota_{g}\phi}v$, where $\phi\in\mathrm{Aut}(G(m))$ is of the form in Eq. 3: Here $v=_{G(m)}g^{-1}\phi(w)^{-1}guw$. Rearranging gives $gv=_{G(m)}\phi(w)^{-1}(gu)w$, and so this case reduces to solving the twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$ with respect to $\phi$. This is solvable in linear time by Proposition 20.
5. Conjugacy problem in extensions of $G(m)$
@@
{defi}
Let $A\leq\mathrm{Aut}(G)$ for a group $G=\langle X\rangle$. The orbit decidability problem for $A$, denoted OD(A), takes as input two words $u,v\in X^{\ast}$, and decides whether there exists $\phi\in A$ such that $v\sim\phi(u)$. When $m$ is odd, we recall that $|\mathrm{Out}(G(m))|=2$ by Theorem 4, so we can apply a similar technique as [12, Theorem 5.1] to prove Theorem 1.
Proof 5.1 (Proof of Theorem 1).
Let $\psi_{1},\dots\psi_{s}\in\mathrm{Aut}(G(m))$ be given, and consider $A=\langle\psi_{1},\dots\psi_{s}\rangle\leq\mathrm{Aut}(G(m))$. For each $i=1,\dots,s$, compute $g_{i}\in G(m)$ such that $\psi_{i}=\iota_{g_{i}}\phi_{i}$, where $\iota_{g_{i}}\in\mathrm{Inn}(G(m))$ and $\phi_{i}\in[\psi_{i}]$. Given two words $u,v\in X^{\ast}$, we want to decide whether $v\sim\psi(u)$ for some $\psi\in A$.
If $[\psi_{i}]$ is trivial, for every $i$, then $A\leq\mathrm{Inn}(G(m))$, and so the set $\{\psi(u)\mid\psi\in A\}$ is a collection of conjugates of $u$. Here our problem reduces to deciding if $v$ is conjugate to $u$, which is decidable by [13, Proposition 3.1]. Otherwise, the set $\{\psi(u)\mid\psi\in A\}$ is a collection of conjugates of $u$ and $\phi(u)$, where $\phi\in\mathrm{Aut}(G(m))$ as in Eq. 3. We therefore need to decide whether $v$ is conjugate to either $u$ or $\phi(u)$, which is decidable by two applications of [13, Proposition 3.1].
Our final result is then immediate from Theorem 1, Theorem 1 and Theorem 23.
@@ Acknowledgements
The author would like to thank Laura Ciobanu and Fujii Michihiko for helpful discussions. They would also like to thank Andrew Duncan and Alessandro Sisto for their comments, as well as the reviewer for suggestions.
@@ Appendix A Geodesic normal forms
We note that for $\epsilon\in\{\pm 1\}$, then any pairs $\left(y^{\epsilon(k+1)},y^{-\epsilon(k+1)}\right)$ can be rewritten as $\left(y^{-\epsilon k},y^{\epsilon k}\right)$, using the rewrite rule
@@
This covers all cases, and we obtain a geodesic representative in linear time.
@@
Table 2. Geodesic normal forms